# CardioAI: A Multimodal AI-based System to Support Symptom Monitoring and Risk Detection of Cancer Treatment-Induced Cardiotoxicity

### v2

@@ 4. System Prototype
Based on clinicians’ insights and feedback on design suggestions, we refined our initial user interface designs. We proposed a multimodal AI-based system for continuous monitoring cancer patients with anti-cancer treatment in a non-clinical setting (i.e., at home) by integrating advanced technologies across three main components: wearable and smart speaker, the backend, and dashboard with data visualization.
The system architecture of our proposed multimodal AI-based system is shown in Figure 3. Our system integrates a wearable device, an LLM-based voice assistant, data processing modules, and a clinician-facing dashboard. The system is designed to seamlessly collect patient data from the wearable and smart speaker, process it through the backend, and present key data summarization and explainable AI risk scores to clinicians.
@@ 4.1. Wearable and Smart Speaker Hardware
Our findings in Section 3.2.2 suggest opportunities to use wearables and LLM-based voice assistants to monitor continuous patient physiological data and patient-reported symptoms.
@@
4.2. Backend
The backend architecture stores, processes, and analyzes the data collected by the wearable and smart speaker with the following modules.
@@ 4.2. Backend
We leverage Transformer (Vaswani, 2017) as the backbone to predict cardiotoxicity after assignment to treatment. Static patient information (e.g., demographics), and temporal data (e.g., prior medical events), are passed into the Transformer model to calculate the probability of cardiotoxicity risk. All learnable parameters, including the scale and shape parameters of the Weibull distribution (Barrett, 2014), the hazard weights, and Transformer parameters, are optimized through a mini-batch stochastic gradient descent process (eon Bottou, 1998). To enhance model interpretability, we adopt the Shapley value method (Winter, 2002), providing clinicians with clear insights into the key factors driving high or low cardiotoxicity risk predictions. Additionally, we incorporate an LLM-based explanation generation module to provide plain-language explanations of risk scores and Shapley values, making the results more accessible to clinicians.
@@
The Summarization Module generates comprehensive daily summaries by retrieving and synthesizing data from Information Database: patient EHRs, previous conversation logs, physiological data. This module, powered by GPT-4 (OpenAI, [n. d.]), ensures that healthcare providers receive a clear, concise, and holistic view of the patient’s health status each day. By pulling data from the Information Database, the module reviews historical EHRs to provide essential context, the most recent conversational data, data from the wearables with a particular focus of detecting any abnormalities or deviations in physiological metrics such as heart rate, respiration, or SpO2 levels. The summarization process highlights key trends and outliers, such as abnormal physiological data or significant changes in reported symptoms, while providing a comprehensive background drawn from the patient’s EHR.
@@ 4.3. Frontend: Information Dashboard UI
We provided an interactive dashboard for clinicians which consolidates and visualizes patient data from the other modules in the system. Based on clinicians’ suggested design considerations from participatory design sessions in Section 3.2.3, we revised our initial user interface design accordingly. This dashboard interface has five main components as shown in Figure 1.
Sec. A: Patient Information Overview integrates seamlessly with existing EHR systems, providing immediate access to critical patient information without the need to switch between different platforms, such as demographics, cancer type, stage, and treatment history.
Sec. B: Daily Summary aggregate and summarize key information from collected self-reported symptoms and wearable sensor data in a concise format to support a quick overview of patient’s current status, particularly focusing on any significant changes or issues that might require immediate intervention.
Sec. C: Wearable Sensor Data displays the collected physiological data, such as heart rate, respiration, oxygen saturation, and skin temperature. The visualization is presented in an interactive format, allowing providers to explore different data modalities over specific time periods, quickly identifying trends or anomalies in the patient’s physiological signals.
Sec. D: AI Risk Prediction presents the predicted cardiotoxicity risk score and its trends. We also present its Shapley Value as a breakdown of the most influential factors that predicted the risk score, and an LLM-generated plain language explanation of the risk score and Shapley Value to enhance explainability.
Sec. E: Conversational Log shows the raw conversation log between patients and the smart speaker. The visualization leverages a clear color scheme to differentiate ”normal” (green) and ”abnormal” (yellow) symptoms reporting and a selection feature to enable clinicians to select specific dates to review the detailed conversation logs.
@@ 5. Evaluation
To evaluate our revised user interface described in Section  4.3, we designed an online Figma (Figma, 2024) prototype that participants could interact directly to mimic real-world clinical use. We conducted user studies with four clinicians to further understand how they use the information in different stages of clinical decision-making and their future expectations.
@@ 5.1. Methods: Participants and Procedure
At the beginning of the study, participants received an online link to a Figma prototype. The prototype displayed realistic and synthetic patient data that we had generated specifically for the study. Participants were asked to share their screens and navigate through various tabs and dashboard interface features. During this process, they were encouraged to think aloud, explaining each step they took to interpret patient data and how the information was related to their current diagnostic workflow. They also shared their thoughts on the interface design.
Following the think-aloud session, we conducted a semi-structured interview with each participant. These interviews explored how they envisioned using the data for decision-making, their impressions of the usefulness and usability in a clinical setting, and their expectations or concerns about integrating the system into their workflow. The study concluded with a brief evaluation questionnaire, including the NASA Task Load Index (TLX) (Hart, 1988) and the System Usability Scale (SUS) (Brooke, 1995), to assess perceived workload and system usability. The TLX (Hart, 1988)questionnaire used a 7-point Likert scale to assess different elements of the task load. For questions related to demand, effort, or frustration, lower scores reflected a lighter workload, which was preferred for the system. For performance questions, higher scores indicated better results. The SUS (Brooke, 1995), a widely recognized usability scale, asked users to rate their agreement with 10 statements, ranging from ”strongly disagree” to ”strongly agree”, on topics like ease of use of the system, the need for technical support, and learning difficulty. Two researchers processed the interview transcriptions with thematic analysis, as in the previous participatory design sessions. We summarize our findings below.
@@ 5.2.1. Clinicians’ Uses of Information for Decision-making
During think-aloud sessions, clinicians described how they would use the information in their decision-making, and we identified how the system could support them in the three stages of the current decision-making workflow as described in our previous study (Wu et al., 2024): symptom identification, diagnostic testing and collaboration with other specialties, and clinical decision-making and intervention.
@@
Clinicians appreciate that the system allows them to assess cardiotoxicity risk before proceeding with treatment. P10 highlighted how this feature supports informed decision-making, particularly when choosing between treatments with varying levels of risk:
@@ 5.2.2. Usability of Our System
During the study, clinicians were able to interpret the presented information and incorporate them into their existing workflow. As they interacted with the system, clinicians highlighted several key aspects of the system’s usability and its potential to support their clinical decision-making. The average SUS score is $72.33\pm 1.89$ (out of 100), suggesting it offers a solid level of usability.
@@
A major theme highlighted by participants was the system’s simplicity and ease of use, which they considered essential for smooth integration into their clincial workflow. All participants rated the system very highly for ease of use in SUS, indicating that they found the system very easy to navigate, such as P3’s comment: “like this interface so far, is relatively simple… It doesn’t look cumbersome.” Similarly, P2 echoed these sentiments, noting that the intuitive design was an asset: “The design is good and simple. I think that’s a great way to start. ” This simplicity in design was further supported by the NASA-TLX effort score, which averaged $1.13\pm 0.70$ (out of 7.0), indicating the system required minimal effort to operate and could be adopted easily.
Participants also expressed high confidence in their ability to use the system effectively, supported by the SUS confidence metric, and many stated that they felt comfortable completing tasks without need for additional help. They agreed that most people would be able to learn to use the system very quickly, a sentiment captured by the high SUS score for this question, highlighting its low learning curve and easy fit within their existing workflows.
@@
Participants emphasized that the system’s ease of access to relevant information was a significant advantage. By presenting the data in a streamlined manner, the system minimized the need to manually search through multiple records, improving workflow efficiency. As P4 explained, “If you have more information, this will make it easier without us having to look back in the chart.” The NASA-TLX results further highlighted that the system did not impose a high workload on users as mental demand ($1.13\pm 0.23$) and physical demand ($0.73\pm 0.42$) are low. This suggests that the system effectively supported the retrieval and organization of information without imposing an additional burden on users’ mental or physical resources.
Participants also appreciated the system’s ability to deliver  highly specific and organized information, particularly in the context of cardiotoxicity-related decision-making. This targeted presentation of data was seen as essential for supporting them to make informed decisions. P4 pointed out that the system effectively prioritized cardiotoxicity-related information with a more focused view:
@@
The system’s capacity to  facilitate proactive decision making was a significant benefit highlighted by the participants. By providing real-time updates and predictive insights, the system allowed clinicians to anticipate potential complications and intervene early, shifting from a reactive to a more proactive approach. P5 remarked,
@@
In addition to enhancing proactive care, the system was praised for its role in  complementing existing clinical tools rather than replacing them. Clinicians felt that the system could augment their capabilities to monitor patients. Clinicians also recognized the AI-based system’s potential for  supporting long-term monitoring and follow-up care, particularly for patients in the survivorship phase of cancer treatment. The system’s ability to continuously monitor patient health for extended periods was seen as a valuable tool to detect later-emerging cardiotoxic effects and ensure ongoing patient safety. P10 suggested,
@@ 5.2.3. Future Expectations
While clinicians highlighted the system’s improved access to relevant information as a significant advantage, they also expressed the need for more clinically relevant data to further support decision-making. One key request was for comprehensive tracking of all patient medications, including non-cancer treatments that could interact with cancer therapies or exacerbate cardiotoxicity. P7 pointed out, “This is only the cancer treatment. Normally, it’s not just 3 or 4 medications… there are other medications,” highlighting the complexity of managing patients undergoing cancer treatment and the necessity for the system to provide a holistic view of all medications and potential interactions. By incorporating a more thorough medication history, the system could better support clinicians in recognizing potential drug interactions and side effects that could impact patient health outcomes.
Clinicians also expressed the need for more detailed information about patients’ anti-cancer treatments, such as the type of chemotherapy administered, its last administration date, cumulative dose, and associated cardiotoxic risks. They emphasized that not all anti-cancer treatments have the same cardiotoxic potential, and understanding these specifics is essential for effective management. P7 remarked, “I would want to know what exactly is the chemotherapy that we’re talking about, because not all chemotherapies have cardiotoxicity,” pointing to the necessity of providing granular details to assess risks accurately and make informed treatment decisions.
Furthermore, while clinicians praised the system’s ability to streamline access to relevant cardiotoxicity-related information, reducing the need to search through multiple records, they also expressed the expectation for it to integrate more seamlessly with the tools they already use. Reflecting on past experiences, they described the challenge of navigating between various systems and managing the multiple risk scores and alerts already in EHRs. As P1 noted,
@@
Moreover, the collaborative nature of decision-making often necessitates input from both cardiologists and oncologists, making it crucial for the system to facilitate information sharing between specialties. P11 highlighted this collaborative dynamic, emphasizing the importance of reviewing notes from other specialists to gain a comprehensive understanding of the patient’s treatment and condition:

### v3

@@ 4. System Prototype
The opportunities and suggested design revisions outlined in Section  3.3.2 guided the development of our pilot CardioAI system, designed to address the four key challenges in managing cancer treatment-induced cardiotoxicity. The system employs a multi-modal AI-based approach, offering continuous remote monitoring and explainable risk prediction to support clinicians’ decision-making. An overview of our pilot system architecture is shown in Figure  5. The pilot system is designed to seamlessly collect patient data from wearable devices and an LLM-VA (Section  4.1), process it through the backend infrastructure (Section  4.2), and present collected information and explainable AI risk scores on a clinician-facing dashboard (Section  4.3).
@@ 4.1. Wearable and Smart Speaker Hardware
Our findings in Section  3.3.2 suggest opportunities (Opportunity 1, 2 & 3) to use wearables and LLM-VA to monitor patient physiological data and patient-reported symptoms continuously.
@@
4.2. Backend Algorithms
The backend architecture stores, processes, and analyzes the data the wearable and smart speaker collects with the following modules.
@@ 4.2. Backend Algorithms
Figure 4. AI-based Cardiotoxicity Risk Score Prediction Framework.
[…]
The figure illustrates the architecture of the cardiotoxicity prediction framework. The framework processes static and temporal patient data to predict cardiotoxicity probability. Static information, such as demographics and risk factors, undergoes linear projection. Temporal information, including observations and collection times, is processed through a variable attention (VA) mechanism and a time embedding layer. The outputs from these layers feed into a transformer block, which comprises multi-head self-attention and a feed-forward neural network. The final output of the transformer block is combined with a linear projection layer and a Weibull Cox proportional hazards (WCPH) model to generate the cardiotoxicity probability.
@@
We leverage Transformer (Vaswani, 2017) as the backbone to predict cardiotoxicity after assignment to treatment. A logistic regression model is used first to analyze long-term EHRs to identify high-risk factors with positive coefficients. These identified risk factors, alongside static patient information (e.g., demographics), and temporal data (e.g., prior medical events), are passed into the Transformer model to calculate the probability of cardiotoxicity risk, as Figure 4 shows. The Transformer model processes sequences of patient visits and generates health state vectors, embedding medical codes, procedures, and medications into fixed-size low-dimensional vectors to mitigate high-dimensionality issues common in EHR data. To focus on the most relevant medical events within a visit, we integrate a variable attention module, which automatically prioritizes critical events for accurate cardiotoxicity prediction. For time-to-event prediction, we deploy a Weibull Cox proportional hazards (WCPH) model, converting the cardiotoxicity risk prediction into a survival analysis framework. The risk score is calculated as a product of a baseline hazard function and the patient-specific cardiotoxicity risk based on the Transformer’s output, reflecting the clinical trajectory of patients over time. All learnable parameters, including the scale and shape parameters of the Weibull distribution (Barrett, 2014), the hazard weights, and Transformer parameters, are optimized through a mini-batch stochastic gradient descent process (eon Bottou, 1998). To improve the interpretability of the model, we adopt the Shapley value method (Winter, 2002), providing clinicians with clear information on the key factors driving the predictions of high or low cardiotoxicity risk. Additionally, we incorporate an LLM-based explanation generation module to provide plain-language explanations of risk scores and Shapley values, making the results more accessible to clinicians.
@@
The Summarization Module generates comprehensive daily summaries by retrieving and synthesizing data from Information Database: patient EHRs, previous conversation logs, and physiological data. This module, powered by GPT-4 (OpenAI, [n. d.]), ensures that healthcare providers receive a clear, concise, and holistic view of the patient’s health status each day. By pulling data from the Information Database, the module reviews historical EHRs to provide essential context, the most recent conversational data, and data from the wearables with a particular focus on detecting any abnormalities or deviations in physiological metrics such as heart rate, respiration, or SpO2 levels.
[…]
This figure illustrates the system architecture for a cardiotoxicity monitoring solution that integrates multimodal data, such as wearable sensor data and voice input, with AI-powered modules for summarization, risk prediction, and conversation. On the left side of the diagram, the patient interacts with the system by providing voice input through a wearable device and smart speaker. The voice input and wearable data are sent to the system, which consists of three primary modules. The first module is the Conversation Module, which uses a large language model (GPT-4) and a Retrieval Augmented Generation (RAG) component that retrieves verified external knowledge (such as NIH guidelines). The retrieved data is used to generate responses for the patient. The second module is the Summarization Module, also powered by GPT-4, which summarizes key information from the data. The third module is the Risk Prediction Module, which uses a Transformer model and Shapley values to predict cardiotoxicity risk and explain the important factors influencing the prediction. All data is stored in an Information Database, and a Dashboard Interface is available to care providers, displaying the summarized data, cardiotoxicity risk scores, feature importance (as determined by Shapley values), and allowing for the inclusion of notes and corrections. The entire system is designed to provide continuous monitoring and decision support for both patients and clinicians, leveraging AI models to assess risk and aid in early diagnosis.
@@ 4.3. Frontend: Information Dashboard UI
Based on the opportunities and design revisions suggested by clinicians in Section 3.3.2, we transformed the initial low-fidelity UI prototype into a detailed and functional final interactive dashboard. While the initial iteration explored multiple design and content options, the final UI reflects an agreed-upon choice that consolidates clinicians’ preferences into a streamlined, user-centered interface, which operationalizes the conceptual representations into practical and cohesive elements tailored to meet clinicians’ workflows and decision-making needs.
Specifically, we reorganized the overall layout to position wearable sensor data and conversation logs on the same page, placing them in parallel, which allows clinicians to view physiological metrics alongside patient-reported symptoms simultaneously without additional navigation. In addition, we prioritized displaying the metrics and symptoms that clinicians identified as most relevant for cardiotoxicity management and adopted visualization formats that clinicians suggested as most appropriate for each metric. To Design Revision 4, we revised the section at the upper-right corner of the dashboard, which defines the cardiotoxicity risk score, includes dynamic visualizations of risk scores over time, and incorporates feature importance indicators to clarify the key factors contributing to the score. Furthermore, we separate the summary of symptoms and metrics from the prior summary module and locate this section below the patient information, which leverages LLM-generated summaries to highlight abnormalities. The final dashboard interface has five main modules, as shown in Figure 1.
Module A: Patient Information Overview integrates seamlessly with existing EHR systems, providing immediate access to critical patient information without the need to switch between different platforms, such as demographics, cancer type, stage, and treatment history.
Module B: Daily Summary aggregate and summarize key information from collected self-reported symptoms and wearable sensor data in a concise format to support a quick overview of the patient’s current status, particularly focusing on any significant changes or issues that might require immediate intervention.
Module C: Wearable Sensor Data displays the collected physiological data, such as heart rate, respiration, oxygen saturation, and skin temperature. The visualization is presented in an interactive format, allowing providers to explore different data modalities over specific time periods, quickly identifying trends or anomalies in the patient’s physiological signals.
Module D: Explainable AI-generated Risk Score presents the predicted cardiotoxicity risk score and its trends. We also present its Shapley Value as a breakdown of the most influential factors that predicted the risk score and an LLM-generated plain-language explanation of the risk score and its Shapley Value.
Module E: Conversation Log shows the raw conversation log between the patients and the smart speaker. The visualization leverages a clear color scheme to differentiate “normal” (green) and “abnormal” (yellow) symptoms reporting and a selection feature to enable clinicians to select specific dates to review the detailed conversation logs.
@@ 5. Evaluation
To understand clinicians’ perceptions of our pilot system, we conducted a heuristic evaluation study with clinicians to interact with and provide feedback on our pilot system. Each individual evaluation focused specifically on assessing the clinician-facing dashboard (Section  4.3) to systematically examine how clinicians would utilize the system for decision-making for cancer treatment-induced cardiotoxicity, further addressing RQ2. The evaluation results demonstrate that participants generally appreciated CardioAI’s simplicity, ease of use, improved access to relevant information, and support for proactive clinical decision-making.
@@ 5.1. Methods: Participants and Procedure
As we focus on evaluating how the pilot system would support clinicians’ decision-making at this phase, our heuristic evaluation served as a design probe to explore specific aspects of the clinician-facing dashboard, such as the presentation of information, accessibility, clinicians’ information usage, and their overall perception. The goal was to refine the system’s design by evaluating its usability and identifying areas for improvement, leveraging targeted insights from a smaller and highly specific group of clinicians before engaging a larger cohort of clinical experts in future evaluations. We adopted the same methodology described in Section  3.1 to recruit clinicians from our previous participatory design study cohort. Through convenience sampling, we recruited four clinicians (P3, P5, P7, P11): two oncologists and two cardiologists. This balanced participant group was selected to ensure high sample specificity (Malterud et al., 2016), representing the primary specialties involved in cancer treatment-induced cardiotoxicity management. Given the study’s narrow aim of evaluating the clinician-facing dashboard’s usability and decision-making support, this specific and well-defined group provided sufficient information power (Malterud et al., 2016) to address our research objectives. Before the study, we pre-loaded synthetic patient data into the prototype, representing cardiotoxicity scenarios typically encountered in clinical practice. At the beginning of each session, researchers first introduced the purpose of the study and provided a brief explanation of the think-aloud methodology. Clinicians then engaged in unstructured interaction with the dashboard. They freely navigated its features and explored its functionalities while verbalizing their thought processes and actions. To observe natural interaction patterns, researchers refrained from providing specific guidance during this phase.
Following each think-aloud session, we then conducted a semi-structured interview with each participant. The interview questions can be found in Appendix  B. These interviews further explored how they envisioned using the dashboard for decision-making, their perceptions of the usability in a clinical setting, and their expectations or concerns about integrating the system into their workflow. The study concluded with a brief evaluation questionnaire including the NASA Task Load Index (TLX) (Hart, 1988) and the System Usability Scale (SUS) (Brooke, 1995), to assess perceived workload and system usability. The TLX (Hart, 1988)questionnaire used a 7-point Likert scale to assess different elements of the task load. For questions related to demand, effort, or frustration, lower scores reflected a lighter workload, which was preferred for the system. For performance questions, higher scores indicated better results. The SUS (Brooke, 1995), a widely recognized usability scale, asked users to rate their agreement with 10 statements, ranging from ”strongly disagree” to ”strongly agree”, on topics like ease of use of the system, the need for technical support, and learning difficulty. Two researchers processed the interview transcriptions with thematic analysis, as in the previous participatory design sessions. We summarize our findings below.
@@ 5.2.1. Clinicians’ Uses of Information for Decision-making
During think-aloud sessions, clinicians described how they would use the information in their decision-making, and we identified how the system could support them in the three stages of the current decision-making workflow as discussed in related work: symptom identification, diagnostic testing and collaboration with other specialties, and clinical decision-making and intervention.
@@
Clinicians appreciate that the system allows them to assess cardiotoxicity risk before proceeding with treatment. P3 highlighted how this feature supports informed decision-making, particularly when choosing between treatments with varying levels of risk:
@@ 5.2.2. Usability of Our System
During the study, clinicians were able to interpret the presented information and incorporate it into their existing workflow. As they interacted with the system, clinicians highlighted several key aspects of the system’s usability and its potential to support their clinical decision-making. The average SUS score is $72.33\pm 1.89$ (out of 100), suggesting it offers a solid level of usability.
@@
A major theme highlighted by participants was the system’s simplicity and ease of use, which they considered essential for smooth integration into their clinical workflow. All participants rated the system very highly for ease of use in SUS, indicating that they found the system very easy to navigate, such as P3’s comment: “like this interface so far, is relatively simple… It doesn’t look cumbersome.” This simplicity in design was further supported by the NASA-TLX effort score, which averaged $1.13\pm 0.70$ (out of 7.0), indicating the system required minimal effort to operate and could be adopted easily.
Participants also expressed high confidence in their ability to use the system effectively, supported by the SUS confidence metric, and many stated that they felt comfortable completing tasks without the need for additional help. They agreed that most people would be able to learn to use the system very quickly, a sentiment captured by the high SUS score for this question, highlighting its low learning curve and easy fit within their existing workflows.
@@
Participants emphasized that the system’s ease of access to relevant information was a significant advantage. By presenting the data in a streamlined manner, the system minimized the need to search through multiple records manually, improving workflow efficiency. As P11 explained, “If you have more information, this will make it easier without us having to look back in the chart.” Furthermore, the NASA-TLX results highlighted that the system did not impose a high workload on users as mental demand ($1.13\pm 0.23$) and physical demand ($0.73\pm 0.42$) are low. This suggests that the system effectively supported the retrieval and organization of information without imposing an additional burden on users’ mental or physical resources.
Participants also appreciated the system’s ability to deliver  highly specific and organized information. This targeted presentation of data was considered essential to help them make informed decisions related to cardiotoxicity. P11 pointed out that the system effectively prioritized cardiotoxicity-related information with a more focused view:
@@
The system’s capacity to  facilitate proactive decision-making was a significant benefit highlighted by the participants. By providing real-time updates and predictive insights, the system allowed clinicians to anticipate potential complications and intervene early, shifting from a reactive to a more proactive approach. P5 remarked,
@@
In addition to enhancing proactive care, the system was praised for its role in  complementing existing clinical tools rather than replacing them. Clinicians felt that the system could augment their capabilities to monitor patients.
Clinicians also recognized the AI-based system’s potential for  supporting long-term monitoring and follow-up care, particularly for patients in the survivorship phase of cancer treatment. The system’s ability to continuously monitor patient health for extended periods was seen as a valuable tool to detect later-emerging cardiotoxic effects and ensure ongoing patient safety. P3 suggested,
@@ 5.2.3. Future Expectations
While clinicians highlighted the system’s improved access to relevant information as a significant advantage, they also expressed the need for more clinically relevant data to support decision-making further. One key request was for comprehensive tracking of all patient medications, including non-cancer treatments that could interact with cancer therapies or exacerbate cardiotoxicity. P7 pointed out, “This is only the cancer treatment. Normally, it’s not just 3 or 4 medications… there are other medications,” highlighting the complexity of managing patients undergoing cancer treatment. By incorporating a more thorough medication history, the system could better support clinicians in recognizing potential drug interactions and side effects that could impact patient health outcomes.
Clinicians also expressed the need for more detailed information about patients’ anti-cancer treatments, such as the type of chemotherapy administered, its last administration date, cumulative dose, and associated cardiotoxic risks. They emphasized that not all anti-cancer treatments have the same cardiotoxic potential, and understanding these specifics is essential for effective management. For example, P7 remarked, “I would want to know what exactly is the chemotherapy that we’re talking about, because not all chemotherapies have cardiotoxicity,” pointing to the necessity of providing granular details to assess risks accurately and make informed treatment decisions.
Furthermore, while clinicians praised the system’s ability to streamline access to relevant cardiotoxicity-related information, reducing the need to search through multiple records, they also expressed the expectation for it to integrate more seamlessly with the tools they already use. Reflecting on past experiences, they described the challenge of navigating between various systems and managing the multiple risk scores and alerts already in EHRs. As P5 noted,
@@
In addition, the collaborative nature of decision-making often requires the input of both cardiologists and oncologists, making it crucial for the system to facilitate information sharing between specialties. P11 highlighted this collaborative dynamic, emphasizing the importance of reviewing notes from other specialists to gain a comprehensive understanding of the patient’s condition: